How would i solve| (x-2)(x²+8x-1).                    (X+4)(x²-5x+9), I have to find the answer to these, then add them together.


1
Expert's answer
2022-01-10T06:07:09-0500
"(x-2)(x^2+8x-1)=0"

"x-2=0 \text{ or }x^2+8x-1=0"

"x-2=0=>x_1=2"

"x^2+8x-1=0"

"x^2+8x+16=17"

"(x+4)^2=17"

"x_2=-4-\sqrt{17}, x_3=-4+\sqrt{17}"


"\{-4-\sqrt{17},-4+\sqrt{17}, 2\}"


"(x+4)(x^2-5x+9)=0"

"x+4=0 \text{ or }x^2-5x+9=0"

"x+4=0=>x_4=-4"

"x^2-5x+9=0"

"D=(-5)^2-4(1)(9)=-11<0"

There are no real solutions.


"\{-4\}"

"-4-\sqrt{17}+(-4+\sqrt{17})+2+(-4)=-10"